OpenStudy (anonymous):

Indefinite integral problem with trig functions using substitution: integral of 5 * sec^2(5x) * tan^5(5x) dx

OpenStudy (anonymous):

u = tan(5x), du = 5(sec(5x))^2 dx

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle 5\int\limits_{~}^{~}\sec^2(5x)\times \tan^5(5x)~dx}\) this?

OpenStudy (solomonzelman):

then yes, u=tan(5x) as was suggested, because the derivative of tan(5x) is 5sec^2(5x) (sitting inside there)

OpenStudy (anonymous):

yes though the five is inside the integral, though that shouldn't matter

OpenStudy (solomonzelman):

yes, so your derivative of "u" is sitting inside the integral for you already

OpenStudy (anonymous):

I see how that works. What am I left with after I use that tan(5x)? tan^4(5x)?

OpenStudy (solomonzelman):

no

OpenStudy (anonymous):

u^5

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{~}^{~}5\sec^2x\tan^5(5x)~dx}\) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}5(\sec~x)^2~(\tan~5x)^5~dx}\) \(\large\color{blue}{\displaystyle u=\tan ~5x}\) \(\large\color{green}{\displaystyle du=5(\sec ~5x)^2~dx}\) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\color{green}{5(\sec~x)^2}~(\color{blue}{\tan~5x})^5~\color{green}{dx} }\) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}(\color{blue}{u})^5~\color{green}{du} }\)

OpenStudy (solomonzelman):

and the \(\large\color{green}{ 5(\sec~5x)^2 }\) went away with \(\large\color{green}{ dx }\) TO BE REPLACED BY \(\large\color{green}{ du }\)

OpenStudy (anonymous):

Oh Okay that makes sense! I always forget the laws for trig functions. Okay can I finish this and see if we got the same answer?

OpenStudy (solomonzelman):

yes, apply the power rule, and then sub the tan(5x) back for "u" as you are done.

OpenStudy (anonymous):

So it should be 1/6 * tan^6(5x)?

OpenStudy (solomonzelman):

yes, with one more thing.... (you left something out)

OpenStudy (solomonzelman):

+ ... ?

OpenStudy (anonymous):

+ C

OpenStudy (solomonzelman):

Yup

OpenStudy (anonymous):

uggggggghhhhh I always forget that one after a long problem

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{\tan^6(5x)}{6} ~+C }\)

OpenStudy (solomonzelman):

I would take almost all points off for leaving out +C

OpenStudy (solomonzelman):

because when you don't write +C you only include 1 possible answer out of the infinite number of answers there are.

OpenStudy (solomonzelman):

Good integration, but +C is very important, because not only \(\large\color{black}{ \displaystyle \frac{\tan^6(5x)}{6} ~ }\), but also ANY \(\large\color{black}{ \displaystyle \frac{\tan^6(5x)}{6} ~+C }\) will have a slope of \(\large\color{black}{ \displaystyle 5\sec^2(5x)\tan^5(5x) }\).

OpenStudy (solomonzelman):

(good job though...) yw

OpenStudy (anonymous):

Thanks Solomon. Your name fits you!

OpenStudy (solomonzelman):

You mean it as the King Solomon? Tnx for the complement, I will take it:D

OpenStudy (anonymous):

absolutely!

OpenStudy (solomonzelman):

You too will surely get there... you are moving very quickly in your math latter.